OpenStudy (anonymous):

what is a three step equation in which the answer is 23?

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

i'd like you to try this on your own do something like this 3x + a = b if x is to be 23 then 3x has to be 3*23 = 69 so make b = a number a little higher than 69 and a will be that amount which will make the right side = 69 eg if you add 5 to 69 then a will be 5 ok - have a try at that

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

ok can solve this 3x + 9 = 78

OpenStudy (anonymous):

One step would be x + 3 = 26 Two steps would be (multiply everything by two) 2x + 6 = 52 Three steps??

OpenStudy (anonymous):

ok what is a three step equation where the answer is 24?

OpenStudy (anonymous):

seperate the x's x + x + 6 = 52